# Late-Time Infrared Cooling in Magnetar-Driven Supernovae

### v1

@@ 4 Discussion and Conclusion
This work examined the cooling power from the radiative transfer simulations of magnetar-driven supernovae presented in Omand2023, focusing on the characterizing the infrared cooling to better optimize future JWST observations. Models show little infrared emission at 1 year, a variety of infrared emission profiles at 3 years, and are dominated by infrared emission at 6 years. The fraction of cooling in the infrared show a sharp dependence on the luminosity of the PWN, being optically-dominated above 1041 erg s-1, infrared-dominated below 1039 erg s-1, and showing a variety of properties at intermediate values. The fraction of cooling in infrared also increases at higher ejecta masses and lower average energy in the PWN spectrum, although this dependence is not as strong. The cooling luminosity of [Ne II] 12.81m increases with increasing PWN luminosity and increasing ejecta mass. Cooling at 1 year is mostly done through optical O and S lines, with some contribution from Mg and Ca. Optical coolants are the same at 3 years, but infrared cooling is more pronounced, with contributions from Ar, Ni, and Ne lines. The most ubiquitous line and strongest infrared coolant at both 3 and 6 years is the [Ne II] 12.81m line. At 6 years, the only strong optical coolant is the [Ca II] 7291.5, 7323.9 doublet; while the mid-infrared has the same coolants as at 3 years, and strong S, Fe, Si, and O lines appear in the far-infrared. For the models with the brightest [Ne II] 12.81m line, the line should be detectable out to with 10 ks exposure time. Supernovae with high inferred magnetic field strength should transition from optical- to infrared-dominated at an earlier timescale, and infrared-dominated supernovae are good candidates for radio follow-up to detect the PWN.
@@ Acknowledgements
The author thanks Jacob Wise, Steve Schulze, and Gavin Lamb for their helpful discussions. C. M. B. O. acknowledges support from the Royal Society (grant Nos. DHF-R1-221175 and DHF-ERE-221005).

### v2

@@ 5 Conclusion
This work examines the cooling power from the radiative transfer simulations of magnetar-driven supernovae presented in Omand and Jerkstrand (2023), focusing on the characterizing the infrared cooling to better optimize future JWST observations. Models show little infrared emission at 1 year, a variety of infrared emission profiles at 3 years, and are dominated by infrared emission at 6 years. The fraction of cooling in the infrared show a sharp dependence on the luminosity of the PWN, being optically-dominated above 1041 erg s-1, infrared-dominated below 1039 erg s-1, and showing a variety of properties at intermediate values. The fraction of cooling in infrared also increases at higher ejecta masses and lower average energy in the PWN spectrum, although this dependence is not as strong. The cooling luminosity of [Ne II] 12.81m increases with increasing PWN luminosity and increasing ejecta mass. Cooling at 1 year is mostly done through optical O and S lines, with some contribution from Mg and Ca. Optical coolants are the same at 3 years, but infrared cooling is more pronounced, with contributions from Ar, Ni, and Ne lines. The most ubiquitous line and strongest infrared coolant at both 3 and 6 years is the [Ne II] 12.81m line. At 6 years, the only strong optical coolant is the [Ca II] 7291.5, 7323.9 doublet; while the mid-infrared has the same coolants as at 3 years, and strong S, Fe, Si, and O lines appear in the far-infrared. For the models with the brightest [Ne II] 12.81m line, the line should be detectable out to with 10 ks exposure time. Supernovae with high inferred magnetic field strength should transition from optical- to infrared-dominated at an earlier timescale, and infrared-dominated supernovae are good candidates for radio follow-up to detect the PWN.
@@ Acknowledgements
The author thanks Jacob Wise, Steve Schulze, and Gavin Lamb for their helpful discussions, and the anonymous referee for their helpful comments. C. M. B. O. acknowledges support from the Royal Society (grant Nos. DHF-R1-221175 and DHF-ERE-221005).